the length of the flower garden is 4ft less than 5 times the width. the area of the proposed garden is 57ft what is the length of the flower garden
Answer by checkley79(3341)

L=5W-4
A=LW
57=(5W-4)W
57=5W^2-4W
5W^2-4W-57=0
(5W-19)(W+3)=0
5W-19=0
5W=19
W=19/5
W=3.8 ANS.
L=5*3.8-4
L=19-4
L=15 ANS.
PROOF:
57=3.8*15
57=57